# Rate of Repeating Tidal Disruption Events with 5–19 years interval constrained by CRTS and ZTF

### v1

@@ Abstract
Results. We found 2 rTDE candidates, AT 2019azh and AT 2024pvu, with time intervals of 13.2 and 17.1 years, respectively. The peak luminosities of CRTS flares are close to those of ZTF flares. For the CRTS flare of AT 2024pvu, using GALEX UV observations near the peak, we measured a blackbody temperature of $\sim 19500$ K, consistent with TDEs and higher than SNe. Moreover, we estimated the expected number of SNe in the sample to be $\lesssim 0.08$, and hence the probability that both CRTS flares are SNe is only 0.3%. Therefore, the possibility that both CRTS flares are SNe can be ruled out, and it is likely that both are TDEs. Using the two rTDEs, we inferred that the TDE rate is 2–3 orders of magnitude higher than the average over 5–19 years prior to TDE detection. Considering another two rTDEs with intervals of $\sim$2 years in the sample and possible rTDEs missed by CRTS, rTDEs with intervals of <20<20 years may account for 25%–60% of the TDE sample.
Conclusions. Long-interval rTDEs have been overlooked by previous works. Exploring the origin of rTDEs requires taking these long-interval rTDEs into account as well. We prefer to explain rTDEs as repeating partial TDEs, but the possibility of independent TDEs cannot be ruled out and requires future observational tests.
@@ 3.3.2 Repeating partial TDEs
How do stars enter such highly eccentric orbits? A widely accepted channel is that a close stellar binary system is tidally dissociated when approaching the SMBH, with one star escaping and the other being captured into a highly eccentric orbit (Hills 1988). The binary must be tight with separation $a_{\rm bin}\lesssim 10\ R_{\odot}$ to survive stellar scattering in the galactic nucleus (e.g, Hills 1988). The semi-major axis of the captured star is primarily related to $a_{\rm bin}$ (Cufari et al. 2022), which is roughly:
@@
Can the observed high fraction of rTDEs of $\sim 25\%-60\%$ in our sample be explained within this framework? Wide binaries in the galactic nucleus would transform into either single stars or close binaries after scattering with fellow stars. Suppose both single stars and close binary stars encounter SMBH with orbits with $R_{p}\approx 1-2R_{t}$. The former produces a single partial TDE, while the latter produces rpTDEs. Then, the fraction of rpTDE in the observed partial TDEs is:
$f({\rm rpTDE})=\frac{f_{\rm close}\bar{N}}{f_{\rm close}\bar{N}+(1-f_{\rm close})},$ (14)
where $f_{\rm close}$ is the fraction of close binaries, and $\bar{N}$ is the average number of flares caused by each rpTDE. Considering full TDEs with $R_{p}\lesssim R_{t}$, the fraction of rpTDE among all TDEs is even lower. We then discuss the reasonable values for $f_{\rm bin}$ and $\bar{N}$.
Studies of nearby open star clusters have shown a low fraction of close binaries of only $\lesssim 2\%$ with $a_{\rm bin}\lesssim 10R_{\odot}$ (e.g, Patience et al. 2002). The fraction in the galactic nucleus lacks precise observational data. Theoretical studies suggest that gravitational perturbations from the SMBH and flyby stars can convert $\sim 20\%-50\%$ of wide binaries into close binaries (Dodici et al. 2025; Hao-Tse Huang and Lu 2025), corresponding to a $f_{\rm close}$ of $\sim 10\%-30\%$ assuming an initial binary fraction of $\sim 40\%-60\%$ (e.g, Duchêne and Kraus 2013).
We are unable to directly verify the number of flares in our sample through observation at present: Table 3 lists the time of the next flare predicted by the rpTDE model, and the earliest (AT 2020vdq and AT 2022dbl) will be in 2026. Despite this, we do not prefer that rTDEs in our sample are rpTDEs with $\bar{N}$ of several dozens for the following reasons. Firstly, dozens of flares require a fractional mass loss of each pTDE of $\lesssim 5\%$, further requiring $R_{p}/R_{t}$ stuck at the edge where partial TDE can just occur, and the possibility of the entire sample being like this is small. Secondly, the average energy of the flares in our sample is on the order of $10^{51}$ erg (Table 3), which is comparable to the typical value of the ZTF TDE sample (van Velzen et al. 2021; Yao et al. 2023). Stars that can withstand dozens of such disruptions must have larger masses, further reducing the probability. Thirdly, in the case of small fractional mass loss, if the star maintains its structure, leading to a small change in the mass loss, the energy of each flare would be similar, which is inconsistent with the observations of our sample (Table 3, although in line with the situation of ASASSN-14ko). However, if the stellar structure evolves significantly due to tidal heating, the number of observable flares may only be a few (Liu et al. 2025). Finally, for AT 2020vdq and AT 2022dbl, Somalwar et al. (2025) and Lin et al. (2024) found no earlier flares before the ZTF detection, implying that the ZTF detection is the first of the series of flares. This situation would have a too low probability if assuming $\bar{N}$ of several dozens. From both theoretical and observational perspectives, it is reasonable to assume $\bar{N}$ of several for rpTDEs. This means that eRASSt J045650.3–203750 with 5 flares may represent a common situation, and ASASSN-14ko with $\gtrsim 30$ flares may be an isolated case.
If assuming $f_{\rm close}\lesssim 2\%$ and $\bar{N}\sim 5$, the inferred fraction of rpTDEs in partial TDEs is only $\lesssim 10\%$, not enough to explain the high rTDE fraction of $\sim 25\%-60\%$. Assuming $\bar{N}\sim 5$, to account for observations, $f_{\rm close}$ must reach $\sim 10\%-30\%$. Such a high close binary fraction predicts a large number of eclipsing binaries in the galactic nucleus, which can be tested by infrared time-domain surveys observationally.
@@
Another way to increase the rTDE fraction in the Hills channel is to consider a larger $R_{p}$ of the binary’s center of mass: the star is captured with $R_{p}>2R_{t}$ initially, with no partial TDE occurring; after orbiting for a period of time, the star’s $R_{p}$ decreases or $R_{t}$ increases, and partial TDE can occur. The requirement for $R_{p}$ for binary star dissociation is:
$R_{p}\lesssim a_{\rm bin}\left(\frac{M_{\rm BH}}{M_{\rm bin}}\right)^{1/3},$ (15)
where $M_{\rm bin}$ is the total mass of the binary. Assuming $M_{\rm bin}=2M_{\star}$, we have $R_{p}/R_{t}\lesssim 0.8a_{\rm bin}/R_{\star}$. This further requires $R_{p}\lesssim 8R_{t}$ assuming $a_{\rm bin}\lesssim 10R_{\star}$. Assuming a full loss cone regime, which is reasonable to explain a high incidence rate, $R_{p}$ approximately follows a uniform distribution (e.g. Rees 1988). Therefore, allowing a larger $R_{p}$ can boost the fraction of rTDEs predicted by the Hills mechanism by a factor of several. The $R_{t}$ of the captured star can be increased through tidal heating and tidal spin-up (Liu et al. 2025). However, quantitative predictions of how much $R_{t}$ can increase still require more detailed numerical simulations. In addition, the $R_{p}$ of the captured star changes continuously and randomly under the gravitational perturbation of the nuclear cluster on a time scale $t_{J}$ of $\sim 10^{7}-10^{9}$ yr, while $a$ remains almost unchanged (Pan & Lai 2026). Considering only this mechanism, there is approximately a 50% chance that the captured stars will fall into the threshold of $R_{p}<2R_{t}$ and eventually end up with rpTDEs. However, there is a competitive mechanism in the above process, as gravitational wave radiation circularizes the orbit and eventually turns the captured star into an extreme mass ratio inspiral. For $M_{\rm BH}\gtrsim 10^{7}M_{\odot}$, the gravitational-wave circularization prevents the occurrence of rpTDE because its time scale is generally shorter than $t_{J}$ assuming $R_{p}\sim 2-8R_{t}$ (Pan & Lai 2026). Therefore, theoretical research has not yet provided a clear mechanism that can effectively convert the majority of captured stars into rpTDEs.
@@
Besides the Hills mechanism, another channel to generate stars with highly eccentric orbits is the scattering of the S star cluster (Bromley et al. 2012), including internal scattering and scattering with intruding stars. If AT 2019azh and AT 2024pvu are rpTDEs formed in this channel, the scattering position has a distance of $\sim 10^{3}$ AU, coinciding with the inner edge of the S cluster in the Milky Way (Eckart and Genzel 1996). However, the S cluster in the Milky Way is not dense enough to generate a large number of rpTDEs with periods of <20<20 years (Alexander 2017). Recently, Jiang and Pan (2025) proposed a model for interpreting QPE observed years after TDE occurs, involving collision between a pre-existing orbiting star and the newly formed TDE disk. This model suggests exceptionally abundant stars in the vicinity of SMBH, which was explained as a result of recently faded AGN, where stars were captured by the AGN accretion disk (Pan and Yang 2021) or formed in the disk in situ (Fan and Wu 2023). Stars in such rich stellar disks may fall towards the SMBH in highly eccentric orbits through stellar scatters or the gravitational perturbation of a massive object, such as a companion SMBH. The rpTDEs formed in this channel would have semi-major axes of $\sim 10^{3}-10^{4}\ R_{g}$, the size of the AGN outer disk, consistent with the observations of our sample. However, whether this mechanism can explain the observed rTDE rate remains to be studied in detail.
@@ 3.3.3 Other possibilities and future observational tests
In summary, we preferred to interpret rTDEs in the sample using the rpTDE model, while cannot rule out the possibility of independent TDEs. The rpTDE predicts that rTDE is periodic, which requires verification through future observations. Very recently, a new flare was detected in IC 3599 (Grupe et al. 2025), which might belong to a series of TDEs with a period of $\sim 18$ years, along with the two flares previously detected. Follow-up observations of IC 3599 may provide conclusive evidence for the existence of rTDEs with a period of >>10 years. In addition, the independent TDE model predicts that the time intervals of rTDE are uniformly distributed. This could not be tested using the current sample because of the large statistical error at the small sample size, and requires larger rTDE samples in the future.
@@ 4 Summary
We proposed two interpretations for the high rTDE fraction. One is that the two TDEs are independent, while the TDE rate in the galaxies in this sample was boosted by 2–3 orders of magnitude by a pc-scale SMBH companion. The other is that both AT 2019azh and AT 2024pvu are rpTDEs of a star in a highly eccentric ($1-e\sim 10^{-3}$) orbit. We prefer the rpTDE interpretation. However, the channel through which the star enters such an orbit remains unclear. It might be due to the Hills mechanism, or the scattering in an ultra-dense stellar disk left by a faded AGN. We put forward predictions based on the two interpretations, which require future observational tests.

### v2

@@ Abstract
Results. We found 2 rTDE candidates, AT 2019azh and AT 2024pvu, with time intervals of 13.2 and 17.1 years, respectively. The peak luminosities of CRTS flares are close to those of ZTF flares. For the CRTS flare of AT 2024pvu, using GALEX UV observations near the peak, we measured a blackbody temperature of $\sim 19500$ K, consistent with TDEs and higher than SNe. Moreover, we estimated the expected number of SNe in the sample to be $\lesssim 0.08$, and hence the probability that both CRTS flares are SNe is only 0.3%. Therefore, the possibility that both CRTS flares are SNe can be ruled out, and it is likely that both are TDEs. Using the two rTDEs, we inferred that the TDE rate is 2–3 orders of magnitude higher than the average over 5–19 years prior to TDE detection. Considering another two rTDEs with intervals of $\sim$2 years in the sample and possible rTDEs missed by CRTS, rTDEs with intervals of <20<20 years may account for 25%–60% of the TDE sample. We prefer to explain rTDEs as repeating partial TDEs. If so, the high fraction of rTDEs suggests that the observed optical TDE rate has been overestimated. However, the possibility of independent TDEs cannot be ruled out and requires future observational tests.
@@ 3.3.2 Repeating partial TDEs
In the rpTDE model, the series of flares caused by the star should be regarded as a single event. What impact will this have on the optical TDE event rate?
To answer this question, we need first to estimate the average number of flares ($\bar{N}$) produced by each rpTDE. At present, we cannot measure $\bar{N}$ in our sample directly through observations. We consider two possible ranges of $\bar{N}$: several, as in the case of eRASSt J045650.3–203750 (5), and several dozens, as in the case of ASASSN-14ko (>30>30). The former requires a fractional mass loss of $\gtrsim 10\%$, while the latter requires it to be $\lesssim 5\%$. We prefer the former for the following reasons. Firstly, a small fractional mass loss requires that $\beta$ be within a narrow range (Ryu et al. 2020a, b), and the possibility of the entire sample being like this is small. Secondly, the average energy of the flares in our sample is on the order of $10^{51}$ erg (Table 3), which is higher than that of ASASSN-14ko of $\sim 10^{50}$ erg (Huang et al. 2025) and comparable to the typical value of the ZTF TDE sample (van Velzen et al. 2021; Yao et al. 2023). It would be more reasonable to assume $\bar{N}$ of several because $\bar{N}$ of several dozen requires disruptions of large mass stars with low probability. Thirdly, the flares’ energies can vary several times in our sample (Table 3), which does not match the situation in ASASSN-14ko, where the flares’ energies are similar. Finally, for AT 2020vdq and AT 2022dbl, Somalwar et al. (2025) and Lin et al. (2024) found no earlier flares before the ZTF detection, implying that the ZTF detection is the first of the series of flares. This situation would have too low a probability if assuming $\bar{N}$ of several dozen. From both theoretical and observational perspectives, it is reasonable to assume $\bar{N}$ of several for rpTDEs.
An rpTDE is more likely to fall within a time-limited sample than a single TDE. This selection effect can lead to overestimation of the TDE rate and the rpTDE fraction. There are a total of 18 ZTF flares in our sample considering that AT 2020vdq and AT 2022dbl each contributed 2 flares. we assume that $\sim 6-12$ of them were produced by rpTDEs, and the remaining $\sim 6-12$ were produced by single TDEs. By dividing the former by $\bar{N}$, which is assumed to be $\sim 3-10$, we estimated the correction factor for the optical TDE rate to be $\sim 40\%-80\%$ ($\sim 2\times 10^{-5}\ {\rm galaxy^{-1}\ yr^{-1}}$ using the observed value of the ZTF, Yao23), and the fraction of rpTDEs to be $\sim 5\%-40\%$. Taking into account rpTDEs with periods of $\gtrsim 20$ years, the TDE rate would be further reduced, while the fraction of rpTDEs would be higher.
How do stars enter highly eccentric orbits with $1-e\approx 10^{-3}$? A widely accepted channel is that a close stellar binary system is tidally dissociated as it approaches the SMBH, with one star escaping and the other captured into a highly eccentric orbit (Hills 1988). The binary must be tight with separation $a_{\rm bin}\lesssim 10\ R_{\odot}$ to survive stellar scattering in the galactic nucleus (e.g, Hills 1988). The semi-major axis of the captured star is primarily related to $a_{\rm bin}$ (Cufari et al. 2022), which is roughly:
@@
Can the fraction of rpTDEs of $\sim 5\%-40\%$ be explained within this framework? When estimating the proportions of different types of TDEs, we only consider single stars and close binaries because wide binaries in the galactic nucleus would transform into either single stars or close binaries after scattering with fellow stars. We assumed that rpTDE can occur in close binaries with $R_{p}$ that allows partial TDE, whereas single TDE occurs in single stars or in close binaries with $R_{p}$ that allows full TDE. According to the loss cone theory, the event rate of partial TDEs is roughly twice that of full TDEs (Stone et al. 2020), so the fraction of $R_{p}$ that meets the condition for partial TDE is about 2/3. Thus, the observed rpTDE fraction further requires a fraction of close binaries in all stars $f_{\rm close}\sim 7\%-60\%$. This fraction is much higher than that observed in nearby open star clusters of only $\lesssim 2\%$ ($a_{\rm bin}\lesssim 10R_{\odot}$, e.g., Patience et al. 2002). Recent theoretical studies have shown that in galactic nuclei, gravitational perturbations from the SMBH and flyby stars can convert $\sim 20\%-50\%$ of wide binaries into close binaries (Dodici et al. 2025; Hao-Tse Huang and Lu 2025). Assuming an initial binary fraction of $\sim 40\%-60\%$ (e.g, Duchêne and Kraus 2013), these effects may increase $f_{\rm close}$ to $\sim 10\%-30\%$, meeting the requirement to explain rpTDEs using the Hills mechanism. This high fraction further predicts a large number of eclipsing binaries in the galactic nucleus, which can be tested observationally through infrared time-domain surveys.
@@
Another way to boost the rpTDE rate in the Hills channel is to consider a larger $R_{p}$ of the binary’s center of mass: the star is captured with $R_{p}>2R_{t}$ initially, with no partial TDE occurring; after orbiting for a period of time, the star’s $R_{p}$ decreases or $R_{t}$ increases, and partial TDE can occur. The $R_{t}$ of the captured star can increase through tidal heating and tidal spin-up (Liu et al. 2025). In addition, the $R_{p}$ of the captured star can decrease under the gravitational perturbation of the nuclear cluster on a time scale of $\sim 10^{7}-10^{9}$ yr, while $a$ remains almost unchanged (Pan & Lai 2026). However, quantitative predictions of how much $R_{t}$ can increase or $R_{p}$ can decrease still require more detailed numerical simulations.
@@
Besides the Hills mechanism, another channel to generate stars with highly eccentric orbits is the scattering of the S star cluster (Bromley et al. 2012), including internal scattering and scattering with intruding stars. If AT 2019azh and AT 2024pvu are rpTDEs formed in this channel, the scattering position has a distance of $\sim 10^{3}$ AU, coinciding with the inner edge of the S cluster in the Milky Way (Eckart and Genzel 1996). However, the S cluster in the Milky Way is not dense enough to generate a large number of rpTDEs (Alexander 2017). Recently, Jiang and Pan (2025) proposed a model for interpreting QPE observed years after TDE occurs, involving collision between a pre-existing orbiting star and the newly formed TDE disk. This model suggests exceptionally abundant stars in the vicinity of SMBH, which was explained as a result of recently faded AGN, where stars were captured by the AGN accretion disk (Pan and Yang 2021) or formed in the disk in situ (Fan and Wu 2023). Stars in such rich stellar disks may fall towards the SMBH in highly eccentric orbits through stellar scatters or the gravitational perturbation of a massive object, such as a companion SMBH. The rpTDEs formed in this channel would have semi-major axes of $\sim 10^{3}-10^{4}\ R_{g}$, the size of the AGN outer disk, consistent with the observations of our sample. However, whether this mechanism can explain the observed rTDE rate remains to be studied in detail.
@@ 3.3.3 Other possibilities and future observational tests
In summary, we prefer to interpret rTDEs in our sample using the rpTDE model, while cannot rule out the possibility of independent TDEs. The rpTDE model predicts that rTDEs are periodic, which requires verification through future observations. In November 2025, a new flare was detected in IC 3599 (Grupe et al. 2025), which might belong to a series of TDEs with a period of $\sim 18$ years, along with the two flares previously detected. Follow-up observations of IC 3599 may provide conclusive evidence for the existence of rTDEs with a period of >>10 years. In addition, the independent TDE model predicts that the time intervals of rTDE are uniformly distributed. This could not be tested using the current sample because of the large statistical error at the small sample size, and requires larger rTDE samples in the future.
@@ 4 Summary
We proposed two interpretations for the high rTDE fraction. One is that the two TDEs are independent, while the TDE rate in the galaxies in this sample was boosted by 2–3 orders of magnitude by a pc-scale SMBH companion. The other is that both AT 2019azh and AT 2024pvu are rpTDEs of a star in a highly eccentric ($1-e\sim 10^{-3}$) orbit. In this case, the observed optical TDE rate is overestimated. Assuming an average of several flares for each rpTDE, we estimated a correction factor for the rate of $\sim 40\%-80\%$. We prefer the rpTDE interpretation. However, the channel through which the star enters such an orbit remains unclear. It might be due to the Hills mechanism, or the scattering in an ultra-dense stellar disk left by a faded AGN. We put forward predictions based on the two interpretations, which require future observational tests.